OpenStudy (anonymous):

Use the Zero Product Property to solve the equation: -3m(2m – 5) = 0

OpenStudy (anonymous):

The zero product property basically means that if two thing multiply to 0, then one or more of them must be zero. So you have to break up the factors into 2 separate equations and solve both of them. -3m = 0 2m - 5 = 0

OpenStudy (anonymous):

ok so 2m-5 would = -3m ?

OpenStudy (anonymous):

not quite. You set them equal to 0, not each other. Equation 1: -3m = 0 Equation 2: 2m - 5 = 0

OpenStudy (anonymous):

And you solve them both to get 2 answers

OpenStudy (anonymous):

m=0 m=5/2 2m=5

OpenStudy (anonymous):

Sorry I got kicked out and the page wouldn't reload. Yeah the answers are m = 0 and m =5/2